# Coupled Laplacian Eigenmaps for Locally-Aware 3D Rigid Point Cloud Matching

### v1

@@ 7 Coupled Laplacian
  normal-$\triangleright$ Scale factor
@@
Select target cross-connections set $F^{\mathcal{T}}\subset\mathcal{V}^{\mathcal{T}}$ Find sets $F^{\mathcal{S}}\subset\mathcal{V}^{S}$ and $F^{\mathcal{S}}_{M}\subset\mathcal{V}^{S}_{M}$ as in Eq. 3 $L^{C}\leftarrow\text{Coupled Laplacin}(W^{\mathcal{T}},W^{S},F^{\mathcal{T}},F^{\mathcal{S}},F^{\mathcal{S}}_{M})$;
@@
        normal-$\triangleright$ Same side
@@
        normal-$\triangleright$ Opposite side
@@ 8 Spectral-Based BSE
where $x_{\text{max}}=x_{\operatorname*{arg\,max}{f}}$ and $x_{\text{min}}=x_{\operatorname*{arg\,min}{f}}$ are the points corresponding to the maximum and minimum values of the Fiedler vector, respectively. In this way, the input surfaces do not need to lie on the same Euclidean frame in order to have comparable length measurements. Even if Fiedler vectors of different shapes may have opposite sign, the inversion of maximum and minimum does not affect the bone length calculation. Note that, as extension for a more precise length computation, it is also possible to select the $M$ maximum and minimum points of the Fiedler vector and compute $x_{\text{max}}$ and $x_{\text{min}}$ as the barycenters of those points. The ratio between source and target lengths gives a scaling factor, $\alpha$, that we use to scale the target surface in order to match the length of the reference. The mirrored version of the source bone, $\mathcal{V}^{S}_{M}$, is then generated by flipping it around the point cloud second principal component and both versions of the reference are registered to the target with RANSAC algorithm [20]. Interestingly, the adjacency matrix of the two references, before and after registration, is the same as the original one and, therefore, we can avoid its computation multiple times.
@@
After the described pre-processing, the core of the algorithm based on the Coupled Laplacian takes place. The sub-set of target nodes used for cross-connections is stochastically selected and the corresponding points on the two versions of the reference are queried as in Eq. 3. Hence, the three graphs can now be connected and the Coupled Laplacian computed from the whole weighted adjacency matrix. For more clarity, the structure of the coupled graph adjacency matrix is depicted in Fig. 6. Therefore, the first $m$ coupled eigenmaps are computed with a given eigensolver and split into the components of each of the three single shapes, as in Eq. 4. Finally, the Grassman distances, $d_{G}(\cdot,\cdot)$, of the QR normalized aligned eigenmaps restricted to the cross-connection sets are compared in order to predict the target side. The latter is equal to $\mathrm{S}^{S}$ if the Grassman distance between target and reference is lower than the distance between target and mirrored reference, otherwise it will be the opposite body side.
@@ 9 Robustness to Source Variation
Method
Femur
Hip
Fibula
Tibia
Best Hausdorff
89.62
100.0
75.76
86.87
Chamfer
85.85
100.0
86.87
90.91
FPFH [57]
83.96
100.0
83.84
93.94
$\textbf{Ours}_{20}$ 93.40 100.0
89.90
94.95
$\textbf{Ours}_{10}$ 93.40 100.0 91.92
93.94
Worst Hausdorff
51.89
89.74
39.39
51.51
Chamfer
54.72
92.31
35.35
38.38
FPFH [57]
34.34
87.18
34.34
41.41
$\textbf{Ours}_{20}$
49.06
94.87
38.38
39.39
$\textbf{Ours}_{10}$ 57.55
84.62
43.43
45.45
@@ 10 Graphical Comparison
Method
Femur
Tibia
Mean
S normal-$\rightarrow$ S Hausdorff
94.37
97.14
90.48
Chamfer
94.29
88.57
94.88
FPFH [57]
97.14
97.14
97.14
@@
H normal-$\rightarrow$ S Hausdorff
72.22
69.44
70.83
Chamfer
69.44
75.00
72.22
FPFH [57]
77.78
75.00
76.39
$\textbf{Ours}_{20}$
69.44
72.22
70.83
@@
S normal-$\rightarrow$ H Hausdorff
57.94
63.00
60.47
Chamfer
62.62
66.00
64.31
FPFH [57] 63.89
63.00
63.45
$\textbf{Ours}_{20}$
63.55
60.00
61.78
$\textbf{Ours}_{10}$
63.55
66.00 64.78

### v2

@@ 7 Coupled Laplacian
  $\triangleright$ Scale factor
@@
Select target cross-connections set $F^{\mathcal{T}}\subset\mathcal{V}^{\mathcal{T}}$
Find sets $F^{\mathcal{S}}\subset\mathcal{V}^{S}$ and $F^{\mathcal{S}}_{M}\subset\mathcal{V}^{S}_{M}$ as in Eq. 3
$L^{C}\leftarrow\text{Coupled Laplacin}(W^{\mathcal{T}},W^{S},F^{\mathcal{T}},F^{\mathcal{S}},F^{\mathcal{S}}_{M})$;
@@
        $\triangleright$ Same side
@@
        $\triangleright$ Opposite side
@@ 8 Spectral-Based BSE
where $x_{\text{max}}=x_{\operatorname*{arg\,max}{f}}$ and $x_{\text{min}}=x_{\operatorname*{arg\,min}{f}}$ are the points corresponding to the maximum and minimum values of the Fiedler vector, respectively. In this way, the input surfaces do not need to lie on the same Euclidean frame in order to have comparable length measurements. Even if Fiedler vectors of different shapes may have opposite sign, the inversion of maximum and minimum does not affect the bone length calculation. Note that, as extension for a more precise length computation, it is also possible to select the $M$ maximum and minimum points of the Fiedler vector and compute $x_{\text{max}}$ and $x_{\text{min}}$ as the barycenters of those points. The ratio between source and target lengths gives a scaling factor, $\alpha$, that we use to scale the target surface in order to match the length of the reference. The mirrored version of the source bone, $\mathcal{V}^{S}_{M}$, representing its contralateral, is then generated by flipping it around the point cloud second principal component and both versions of the reference are registered to the target with RANSAC algorithm [19]. Interestingly, the adjacency matrix of the two references, before and after registration, is the same as the original one and, therefore, we can avoid its computation multiple times.
@@
After the described pre-processing, the core of the algorithm based on the Coupled Laplacian takes place. The sub-set of target nodes used for cross-connections is stochastically selected and the corresponding points on the two versions of the reference are queried as in Eq. 3. Alternatively, assuming not flat surfaces, other methods for cross-connections can be used instead of NN, e.g. spectral clustering medoids. Hence, the three graphs can now be connected and the Coupled Laplacian computed from the whole weighted adjacency matrix. For more clarity, the structure of the coupled graph adjacency matrix is depicted in Fig. 6. Therefore, the first $m$ coupled eigenmaps are computed with a given eigensolver and split into the components of each of the three single shapes, as in Eq. 4. Finally, the Grassman distances, $d_{G}(\cdot,\cdot)$, of the QR normalized aligned eigenmaps restricted to the cross-connection sets are compared in order to predict the target side. The latter is equal to $\mathrm{S}^{S}$ if the Grassman distance between target and reference is lower than the distance between target and mirrored reference, otherwise it will be the opposite body side.
@@ 9 Robustness to Source Variation
Method Femur Hip Fibula Tibia
Best Hausdorff 89.62 100.0 75.76 86.87
Chamfer 85.85 100.0 86.87 90.91
FPFH [57] 83.96 100.0 83.84 93.94
$\textbf{Ours}_{20}$ 93.40 100.0 89.90 94.95
$\textbf{Ours}_{10}$ 93.40 100.0 91.92 93.94
Worst Hausdorff 51.89 89.74 39.39 51.51
Chamfer 54.72 92.31 35.35 38.38
FPFH [57] 34.34 87.18 34.34 41.41
$\textbf{Ours}_{20}$ 49.06 94.87 38.38 39.39
$\textbf{Ours}_{10}$ 57.55 84.62 43.43 45.45
@@ 10 Graphical Comparison
Method Femur Tibia Mean
S $\rightarrow$ S Hausdorff 94.37 97.14 90.48
Chamfer 94.29 88.57 94.88
FPFH [57] 97.14 97.14 97.14
@@
H $\rightarrow$ S Hausdorff 72.22 69.44 70.83
Chamfer 69.44 75.00 72.22
FPFH [57] 77.78 75.00 76.39
$\textbf{Ours}_{20}$ 69.44 72.22 70.83
@@
S $\rightarrow$ H Hausdorff 57.94 63.00 60.47
Chamfer 62.62 66.00 64.31
FPFH [57] 63.89 63.00 63.45
$\textbf{Ours}_{20}$ 63.55 60.00 61.78
$\textbf{Ours}_{10}$ 63.55 66.00 64.78
@@ 11 Cross-Species BSE
12 Scalability
Figure 15: Average computational time of the eigendecomposition using different combinations of number of shapes ($1$, $3$ or $5$) and modes ($10$ or $20$).
Figure 16: Coupled modes correctly aligned in a partial matching scenario.
The efficiency trade-off of decomposing larger Laplacians is illustrated in Fig. 15. The average computational time is reported for the eigendecomposition, using $10$ and $20$ maps, coupling the same shape with itself $3$ or $5$ times. As expected, larger, i.e. number of shapes, and denser, i.e. number of cross-connections, is the Laplacian, higher is the time needed for the computation. Scalability might be a concern in tasks like shape retrieval with numerous classes. However, simple measures, e.g. downsampling shapes and interpolating eigenmaps, can reduce the complexity. Given the intrinsic coupling of modes, faster, but randomized, solvers, e.g. AMG, are still precise and have been used, cutting the need of resources.
13 Partial Matching
In Fig. 16 we show a graphical example of coupled eigenmaps derived from partially matching shapes. More in detail, an entire femur bone is matched with the distal part, i.e. lower part, of another femur. In this case, artificial cross-connections are added only on the partial match of the source shapes, leading to correctly aligned eigenmaps. Namely, the embeddings are successfully coupled in the distal area of the femur while they are independent on the remaining surface of the source. This example shows the potential of the Coupled Laplacian operator in different, and more complex, scenarios.